# Mean uniformly stable function and its application to almost sure stability analysis of randomly switched time-varying systems

### v2

@@ III-C Almost sure stability of time-varying systems with semi-Markov switching
if $\lambda_{i}(t)$ are MUSFs for all $i\in\mathbb{S}$ and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})<0$.
@@
Then (A.4) yields that
$v_{r_{l+1}}(t_{l+1})\leq\mu_{r_{l+1}r_{l}}v_{r_{l}}(t_{l+1}^{-}).$
Thus, by iteration, there holds
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\int_{0}^{t}\lambda_{r(h)}(h)\mathrm{d}h$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i\in\mathbb{S}}\sum_{k=0}^{N_{i}(t,0)}\int_{t_{i_{k}}}^{t_{i_{k+1}}}\lambda_{i}(h)\mathrm{d}h$
@@
$\leq\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\lim_{t\rightarrow\infty}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\varphi_{i}(S_{i}(k)).$ (22)
@@
$\lim_{t\rightarrow\infty}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\varphi_{i}(S_{i}(k))$ $=\lim_{N_{i}(t,0)\rightarrow\infty}\frac{1}{N_{i}(t,0)}\sum_{k=1}^{N_{i}(t,0)}\varphi_{i}(S_{i}(k))$
$=\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right],\ \mathrm{a.s.}$
Recalling the Lemma 1 and applying strong law of large numbers [34], we can get
(III-C)$=\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right]=\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right],\ \mathrm{a.s.}$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\left(\int_{0}^{t}\lambda_{r(h)}(h)\mathrm{d}h+\sum_{i,j\in\mathbb{S}}N_{ij}(t,0)\ln\mu_{ij}\right)\leq\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}\left(\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right]+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij}\right)<0,\ \mathrm{a.s.}$ (23)
@@
$\ln V_{r}(t,x)\leq\ln V_{r_{0}}(t_{0},\phi)+\sum_{i,j\in\mathbb{S}}N_{ij}(t,t_{0})\ln\mu_{ij}+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h$ (25)
@@
$\limsup_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|\leq$ $\frac{1}{p}\limsup_{t\rightarrow\infty}\frac{1}{t}\left(\ln V_{r_{0}}(t_{0},\phi)+\sum_{i,j\in\mathbb{S}}N_{ij}(t,t_{0})\ln\mu_{ij}+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+\ln c\right)$
@@
Then, together with (S.4), we get (10). Time-varying system with semi-Markov switching is GES a.s. We complete the proof. ∎
@@ III-D Almost sure stability of time-varying systems with Markov switching
A Markov chain is a special form of a semi-Markov chain, and it can be directly derived from a semi-Markov chain by assuming that sojourn time satisfies exponential distributions (see Lemma 2 in [32]). Therefore, we can promptly obtain stability criteria for the system (8) with a Markov switching signal.
@@
Remark 5.
By letting $m_{i}=\frac{1}{q_{i}}$, it is trial to prove the corollaries according to the proof of Theorem 2 and Theorem 2.
@@ III-E Almost sure stability of time-varying systems with renewal process switching
Proof of Theorem 6: Recalling the proof of Theorem 2, condition (A.1-3) yield that
$|x(t)|\leq\alpha_{1}^{-1}\left(\alpha_{2}(|\phi|)e^{\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+N(t,t_{0})\ln\mu}\right).$ (26)
@@
$\int_{t_{k}}^{t_{k+1}}\lambda_{r(t_{k})}(h)\mathrm{d}h=\int_{t_{k}}^{t_{k+1}}\lambda_{i}(h)I_{r(t_{k})=i}\mathrm{d}h\leq\varphi_{i}(S(k))I_{r(t_{k})=i}.$
Since the sequence $\{S(k)\}_{k\in\mathbb{N}^{+}}$ is independent and identically distributed, variables $\varphi_{i}(S(k)),\ k\in\mathbb{N}^{+}$ are also independent and identically distributed, $\mathbb{E}[\varphi_{i}(S(k))]=\mathbb{E}[\varphi_{i}(S(k+1))]$ holds for all $k$ naturally. Define $Z_{k}:=\{\varphi_{i}(S(k))I_{r(t_{k})=i}\}$, we can notice that $Z_{k}$ is an i.i.d. process and
@@
Thus, $\{Z_{k}\},k=0,1,\cdots,N(t,t_{0})$ is a wide-sense stationary process. Meanwhile, we can follow the procedure of Section 3.1 to prove that following equation holds for $l=0,1$:
$\lim_{N(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\int_{t_{k}}^{t_{k+1}}\lambda_{r(t_{k})}(h)\mathrm{d}h$ $\leq\lim_{N(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\varphi_{r(t_{k})}(S(k))$
@@
The time average of $\{Z_{k}\}$ is given by
$A_{z}=\lim_{N_{i}(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\varphi_{i}(S(k))I_{r(t_{k})=i},$
$\bar{A}_{z}=\mathbb{E}[A_{z}]=\lim_{N_{i}(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\mathbb{E}\left[\varphi_{i}(S(k))I_{r(t_{k})=i}\right]=\mathbb{E}[Z_{k}].$
We can verify the Chebychev’s inequality $\mathbb{P}\left[|A_{z}-\bar{A}_{z}|<\varepsilon\right]\geq 1-\frac{\mathrm{Var}(A_{z})}{\varepsilon^{2}}$ holds for some positive $\varepsilon$ with probability 1 by calculating variance of $A_{z}$:
$\mathrm{Var}(A_{z})$ $=\mathbb{E}\left[(A_{z}-\bar{A}_{z})^{2}\right]$
$=\mathbb{E}\left[\lim_{N\rightarrow\infty}(\frac{1}{N+l}\sum_{k=0}^{N+l}Z_{k}-\mathbb{E}[Z_{0}])(\frac{1}{N+l}\sum_{m=0}^{N+l}Z_{m}-\mathbb{E}[Z_{0}])\right]$
$=\lim_{N\rightarrow\infty}\frac{1}{(N+l)^{2}}\sum_{k=0}^{N+l}\sum_{m=0}^{N+l}\mathbb{E}\left[(Z_{k}-\mathbb{E}[Z_{0}])(Z_{m}-\mathbb{E}[Z_{0}])\right]$
$=\lim_{N\rightarrow\infty}\frac{1}{(N+l)^{2}}\sum_{k=0}^{N+l}\sum_{l=0}^{N+l}\left(\mathbb{E}\left[Z_{k}Z_{m}\right]-\mathbb{E}[Z_{0}]^{2}\right)$
$=\lim_{N\rightarrow\infty}\frac{1}{(N+l)^{2}}\sum_{k=0}^{N+l}\sum_{l=0}^{N+l}\left(\mathbb{E}\left[Z_{k}\right]\mathbb{E}\left[Z_{m}\right]-\mathbb{E}[Z_{0}]^{2}\right)$
$=0,$
where $N:=N(t,t_{0})$, the 4th equation follows from the independence property of $\{Z_{k}\}$ and the 5th equation follows from $\mathbb{E}[Z_{k}]=\mathbb{E}[Z_{0}]$. Therefore, (III-E) holds almost surely. By the property of renewal process presented in Lemma 1 and the strong law of large number, we have for $l=0,1$
$\lim_{t\rightarrow\infty}\frac{1}{t}\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h$ $=\lim_{t\rightarrow\infty}\frac{N(t,t_{0})+l}{t}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\int_{t_{k}}^{t_{k+1}}\lambda_{r(t_{k})}(h)\mathrm{d}h$
$\leq\lim_{t\rightarrow\infty}\frac{N(t,t_{0})+l}{t}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\varphi_{i}(S(k))I_{r(t_{k})=i}$
$=\frac{1}{\theta}\sum_{i\in\mathbb{S}}\mathbb{E}\left[\varphi_{i}(S(0))I_{r(t_{0})=i}\right]$
$=\frac{1}{\theta}\sum_{i\in\mathbb{S}}\mathbb{E}\left[\varphi_{i}(S(0))\right]p_{i},\ \mathrm{a.s.}$
and $\lim_{t\rightarrow\infty}\frac{N(t,t_{0})\ln\mu}{t}=\frac{\ln\mu}{\theta},\ \mathrm{a.s.}$
Then combining with (R.4), we have
$\lim_{t\rightarrow\infty}\frac{1}{t}\left(\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+N(t,t_{0})\ln\mu\right)\leq\frac{1}{\theta}\sum_{i\in\mathbb{S}}\mathbb{E}[\varphi_{i}(S(0))]p_{i}+\frac{\ln\mu}{\theta}<0,\ \mathrm{a.s.}$ (28)
which deduces that
$\lim_{t\rightarrow\infty}\left(\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+N(t,t_{0})\ln\mu\right)=-\infty,\ \mathrm{a.s.}$ (29)
We get the same result as (24), so the rest of the proof follows the same steps as the proof of Theorem 2. Here we omit it.
Proof of Theorem 7: Following the proof Theorem 3, (A.2) and (A.3) yield that
$\ln V_{r}(t,x)\leq\ln V_{r_{0}}(t_{0},\phi)+N(t,t_{0})\ln\mu+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h$
Then it follows (B.1) and (28) that
$\limsup_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|\leq$ $\frac{1}{p}\limsup_{t\rightarrow\infty}\frac{1}{t}\left(\ln V_{r_{0}}(t_{0},\phi)+N(t,t_{0})\ln\mu+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+\ln c\right)$
$\leq$ $\frac{1}{\theta p}\left(\sum_{i\in\mathbb{S}}\mathbb{E}\left[\varphi_{i}(S(0))\right]p_{i}+\ln\mu\right)<0,\ \mathrm{a.s.}$
The remaining proof can be completed by following the proof of Theorem 3. ∎
@@ IV Numerical examples
where
$f_{1}(t,x)=\left[\begin{array}[]{c}-2x_{1}+x_{2}\\ x_{1}-2x_{2}\end{array}\right],$
$f_{2}(t,x)=\left[\begin{array}[]{c}-\frac{3}{2}t^{2}x_{1}\\ \cos tx_{1}-\frac{3}{2}t^{2}x_{2}\end{array}\right],$
$f_{3}(t,x)=\left[\begin{array}[]{c}\frac{1}{2}(t\cos t-\frac{1}{2})x_{1}-x_{1}\\ x_{1}-\frac{1}{2}x_{2}+\frac{1}{2}t\left(\cos t-\frac{1}{2}\right)x_{2}\end{array}\right].$
The switching signal $r(t)$ is a semi-Markov chain with a embeded Markov chain determined by transition probability matrix
@@
Choosing Lyapunov functions as follows:
$V_{1}(t,x)=\frac{1}{2}\left(x_{1}^{2}+x_{2}^{2}\right),\ V_{2}(t,x)=x_{1}^{2}+x_{2}^{2},\ V_{3}(t,x)=\frac{1}{2}x_{1}^{2}+x_{2}^{2},$
we thus get $\mu_{21}=\mu_{23}=2$, $\mu_{12}=\mu_{32}=0.5$ and $\mu_{13}=\mu_{31}=1$. By calculating, we get that
$\dot{V}_{1}(t,x)\leq-2V_{1}(t,x),$
$\dot{V}_{2}(t,x)\leq(-3t^{2}+\cos t)V_{3}(t,x),$
$\dot{V}_{3}(t,x)\leq t\left(\cos t-\frac{1}{2}\right)V_{3}(t,x),$
which implies that
$\lambda_{1}(t)=-2,\ \lambda_{2}(t)=-3t^{2}+\cos t,\ \lambda_{3}(t)=t\left(\cos t-\frac{1}{2}\right).$
By applying Theorem 2, we know that system (30) is GUAS a.s. if all the $\lambda_{i}(t)$ are MUSFs and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})<0$. Here, we verify
@@
$\mathbb{E}\left[\int_{t}^{t+S_{3}}\lambda_{3}(h)\mathrm{d}h\right]\leq\mathbb{E}[-\frac{1}{4}S_{3}^{2}+2S_{3}+2]\leq-\frac{1}{4}m_{3}^{2}+2m_{3}+2=5,$
and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})\approx-1.39<0$. We can conclude that system (30) with semi-Markov switching signal is GUAS a.s.
Figure 1 shows the state trajectories of system (30) with semi-Markov switching under single experiment. The dash-dotted line represents the trajectory of $x_{1}$ and the dashed line corresponds to the trajectory of $x_{2}$. It is evident from the figure that the states converge to 00 under a single experiment of semi-Markov switching. Figure 2 and 3 present trajectories under 100 times semi-Markov switching experiments. In all trajectories, the states uniformly asymptotically converge to the origin, confirming the almost sure GUAS property of system (30).
@@
Figure 2: States’ trajectories of system (30) under 100 experiments.
Figure 3: States’ trajectories of system (30) under 100 experiments.
Example 2. Consider the following time-varying system with a Markov switching signal
$\dot{x}(t)=f_{r}(t,x(t))$
where
$f_{1}=\left[\begin{array}[]{c}\frac{1}{2}x_{1}\\ \frac{1}{2}x_{2}\end{array}\right],$
$f_{2}=\left[\begin{array}[]{c}-\frac{1}{2}(3t+1)x_{1}+x_{2}\\ \cos(x_{1})x_{2}-(\frac{3t}{2}+1)x_{2}\end{array}\right],$
$f_{3}=\left[\begin{array}[]{c}\frac{1}{2}(-2+\cos 4t)x_{1}\\ x_{1}+\frac{1}{2}(-3+\cos 4t)x_{2}\end{array}\right].$
Switching signal $r(t)$ is a Markov chain with transition rate matrix
@@
Example 2. Consider the following time-varying system with a Markov switching signal
$\dot{x}(t)=f_{r}(t,x(t))$
where
$f_{1}=\left[\begin{array}[]{c}\frac{1}{2}x_{1}\\ \frac{1}{2}x_{2}\end{array}\right],$
$f_{2}=\left[\begin{array}[]{c}-\frac{1}{2}(3t+1)x_{1}+x_{2}\\ \cos(x_{1})x_{2}-(\frac{3t}{2}+1)x_{2}\end{array}\right],$
$f_{3}=\left[\begin{array}[]{c}\frac{1}{2}(-2+\cos 4t)x_{1}\\ x_{1}+\frac{1}{2}(-3+\cos 4t)x_{2}\end{array}\right].$
Switching signal $r(t)$ is a Markov chain with transition rate matrix
Then, we get from $Q$ that stationary distribution is $\pi=[0.4,0.4,0.2]$.
We choose the Lyapunov function for each subsystem: $V_{1}(t,x)=\frac{1}{2}\left(x_{1}^{2}+x_{2}^{2}\right),\ V_{2}(t,x)=x_{1}^{2}+x_{2}^{2},\ V_{3}(t,x)=\frac{1}{2}x_{1}^{2}+x_{2}^{2}$, and thus get $\mu_{21}=\mu_{23}=2$, $\mu_{12}=\mu_{32}=0.5$ and $\mu_{13}=\mu_{31}=1$. It can check that $\lambda_{1}(t)=1,\ \lambda_{2}(t)=-6t,\ \lambda_{3}(t)=-2+\cos 4t$ and the 1st subsystem is unstable, and the 2nd and 3rd subsystems are stable. Then, we apply the Corollary 4 and verify
$\sum_{i\in\mathbb{S}}\left(\mathbb{E}\left[\varphi_{i}(S_{i})\right]q_{i}+\sum_{j\in\mathbb{S}}q_{ij}\ln\mu_{ij}\right){\pi_{i}}\approx-1.17<0.$
Hence, system (30) with Markov switching signal is GUAS a.s. This example shows that our stability criteria still work for randomly switched systems with unstable subsystems.

### v3

@@ III-C Almost sure stability of time-varying systems with semi-Markov switching
$\lambda_{i}(t)$ are MUSFs for all $i\in\mathbb{S}$ and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})<0$.
@@
Then (A.4) yields that $v_{r_{l+1}}(t_{l+1})\leq\mu_{r_{l+1}r_{l}}v_{r_{l}}(t_{l+1}^{-}).$ Thus, by iteration, there holds
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\int_{0}^{t}\lambda_{r(h)}(h)\mathrm{d}h$
$=$ $\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{i\in\mathbb{S}}\sum_{k=0}^{N_{i}(t,0)}\int_{t_{i_{k}}}^{t_{i_{k+1}}}\lambda_{i}(h)\mathrm{d}h$
@@
$\leq$ $\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\varphi_{i}(S_{i}(k)).$ (20)
@@
$\lim_{t\rightarrow\infty}\frac{1}{N_{i}(t,0)}\sum_{k=0}^{N_{i}(t,0)}\varphi_{i}(S_{i}(k))=\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right],\ \mathrm{a.s.}$
Recalling the Lemma 1 and applying strong law of large numbers [34], following holds almost surely
(III-C)$\!=\!\lim_{t\rightarrow\infty}\sum_{i\in\mathbb{S}}\frac{N_{i}(t,0)}{t}\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right]\!=\!\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right],$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\left(\int_{0}^{t}\lambda_{r(h)}(h)\mathrm{d}h+\sum_{i,j\in\mathbb{S}}N_{ij}(t,0)\ln\mu_{ij}\right)$
$\leq$ $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}\left(\mathbb{E}\left[\varphi_{i}(S_{i}(0))\right]+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij}\right)<0,\ \mathrm{a.s.}$ (21)
@@
$\ln V_{r}(t,x)\leq$ $\ln V_{r_{0}}(t_{0},\phi)+\sum_{i,j\in\mathbb{S}}N_{ij}(t,t_{0})\ln\mu_{ij}$
$+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h$ (23)
@@
$\limsup_{t\rightarrow\infty}\frac{1}{t}\ln|x(t)|$
$\leq$ $\frac{1}{p}\limsup_{t\rightarrow\infty}\frac{1}{t}\left(\ln V_{r_{0}}(t_{0},\phi)\right.$
$\left.+\sum_{i,j\in\mathbb{S}}N_{ij}(t,t_{0})\ln\mu_{ij}+\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h+\ln c\right)$
@@
Then, together with (S.4), we get (9). Time-varying system with semi-Markov switching is GES a.s. ∎
@@ III-D Almost sure stability of time-varying systems with Markov switching
A Markov chain is a special form of a semi-Markov chain, and it can be directly derived from a semi-Markov chain by assuming that sojourn time satisfies exponential distributions (see Lemma 2 in [32]). Therefore, we can promptly obtain stability criteria for the system (8) with a Markov switching signal by letting $m_{i}={1}/{q_{i}}$.
@@ III-E Almost sure stability of time-varying systems with renewal process switching
The proof procedure of Theorems 6 and 7 is very similar to Theorems 2 and 3, we only need to prove:
$\lim_{t\rightarrow\infty}\frac{1}{t}\int_{t_{0}}^{t}\lambda_{r(h)}(h)\mathrm{d}h=\frac{1}{\theta}\sum_{i\in\mathbb{S}}\mathbb{E}\left[\varphi_{i}(S(0))\right]p_{i},\ \mathrm{a.s.}$ (24)
@@
$\int_{t_{k}}^{t_{k\!+\!1}}\!\lambda_{r(t_{k})}(h)\mathrm{d}h\!=\!\int_{t_{k}}^{t_{k\!+\!1}}\!\lambda_{i}(h)I_{r(t_{k})=i}\mathrm{d}h\\ \!\leq\!\varphi_{i}(S(k))I_{r(t_{k})=i}.$
Since the sequence $\{S(k)\}_{k\in\mathbb{N}^{+}}$ is an i.i.d. process, $\left\{\varphi_{i}(S(k))\right\}_{k\in\mathbb{N}^{+}}$ is also an i.i.d. process. $\mathbb{E}[\varphi_{i}(S(k))]=\mathbb{E}[\varphi_{i}(S(k+1))]$ holds for all $k$ naturally. Define $Z_{k}:=\{\varphi_{i}(S(k))I_{r(t_{k})=i}\}$, we can notice that $Z_{k}$ is an i.i.d. process and
@@
Thus, $\{Z_{k}\},k=0,1,\cdots,N(t,t_{0})$ is a wide-sense stationary process and also satisfies ergodicity. Then, for $l=0,1$, there holds
$\lim_{N(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\int_{t_{k}}^{t_{k+1}}\lambda_{r(t_{k})}(h)\mathrm{d}h$
$\leq$ $\lim_{N(t,t_{0})\rightarrow\infty}\frac{1}{N(t,t_{0})+l}\sum_{k=0}^{N(t,t_{0})+l}\varphi_{r(t_{k})}(S(k))$
@@
We can promptly prove (24) by (III-E). This completes the proof.
@@ IV Numerical examples
where $f_{1}(t,x)=[-2x_{1}+x_{2},x_{1}-2x_{2}]^{T}$, $f_{2}(t,x)=[-\frac{3}{2}t^{2}x_{1},\cos tx_{1}-\frac{3}{2}t^{2}x_{2}]^{T}$, $f_{3}(t,x)=[\frac{1}{2}(t\cos t-\frac{1}{2})x_{1}-x_{1},x_{1}-\frac{1}{2}x_{2}+\frac{1}{2}t\left(\cos t-\frac{1}{2}\right)x_{2}]^{T}$. The switching signal $r(t)$ is a semi-Markov chain with a embeded Markov chain determined by transition probability matrix
@@
Choosing Lyapunov functions $V_{1}(t,x)\!=\!\frac{1}{2}\left(x_{1}^{2}\!+\!x_{2}^{2}\right),\ V_{2}(t,x)\!=\!x_{1}^{2}\!+\!x_{2}^{2},\ V_{3}(t,x)\!=\!\frac{1}{2}x_{1}^{2}\!+\!x_{2}^{2},$, we thus get $\mu_{21}=\mu_{23}=2$, $\mu_{12}=\mu_{32}=0.5$ and $\mu_{13}=\mu_{31}=1$. By calculating, we get that $\dot{V}_{1}(t,x)\leq-2V_{1}(t,x),\ \dot{V}_{2}(t,x)\leq(-3t^{2}+\cos t)V_{3}(t,x),\ \dot{V}_{3}(t,x)\leq t\left(\cos t-\frac{1}{2}\right)V_{3}(t,x),$ which implies that $\lambda_{1}(t)=-2,\ \lambda_{2}(t)=-3t^{2}+\cos t,\ \lambda_{3}(t)=t\left(\cos t-\frac{1}{2}\right).$ By applying Theorem 2, we know that system (26) is GUAS a.s. if all the $\lambda_{i}(t)$ are MUSFs and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})<0$. Here, we verify
@@
$\mathbb{E}\left[\int_{t}^{t+S_{3}}\lambda_{3}(h)\mathrm{d}h\right]\leq\mathbb{E}[-\frac{1}{4}S_{3}^{2}+2S_{3}+2]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\leq-\frac{1}{4}m_{3}^{2}+2m_{3}+2=5,$
and $\sum_{i\in\mathbb{S}}\frac{\pi_{i}}{m_{i}}(\mathbb{E}(\varphi_{i}(S_{i}))+\sum_{j\in\mathbb{S}}p_{ij}\ln\mu_{ij})<-1.39$. We can conclude that system (26) with semi-Markov switching signal is GUAS a.s.
Figure 1 shows the state trajectories of system (26) with semi-Markov switching under single experiment. The dash-dotted line represents the trajectory of $x_{1}$ and the dashed line corresponds to the trajectory of $x_{2}$. It is evident from the figure that the states converge to 00 under a single experiment of semi-Markov switching. Figure 2 and 3 present trajectories under 2000 times semi-Markov switching experiments. In all trajectories, the states uniformly asymptotically converge to the origin, confirming the almost sure GUAS property.
@@
Figure 2: States’ trajectories of system (26) under 2000 experiments.
Figure 3: States’ trajectories of system (26) under 2000 experiments.
Example 2. Let’s consider the system (26) with following subsystems and a Markov switching signal: $f_{1}=[\frac{1}{2}x_{1},\frac{1}{2}x_{2}]^{T}$, $f_{2}=[-\frac{1}{2}(3t+1)x_{1}+x_{2},\cos(x_{1})x_{2}-(\frac{3t}{2}+1)x_{2}]^{T}$ and $f_{3}=[\frac{1}{2}(-2+\cos 4t)x_{1},x_{1}+\frac{1}{2}(-3+\cos 4t)x_{2}]^{T}$. Switching signal $r(t)$ is a Markov chain with transition rate matrix
@@
We get from $Q$ that stationary distribution is $\pi\!=\![0.4,0.4,0.2]$.
We choose Lyapunov functions for subsystems: $V_{1}(t,x)=\frac{1}{2}\left(x_{1}^{2}+x_{2}^{2}\right),\ V_{2}(t,x)=x_{1}^{2}+x_{2}^{2},\ V_{3}(t,x)=\frac{1}{2}x_{1}^{2}+x_{2}^{2}$, and thus get $\mu_{21}=\mu_{23}=2$, $\mu_{12}=\mu_{32}=0.5$ and $\mu_{13}=\mu_{31}=1$. It can check that $\lambda_{1}(t)=1,\ \lambda_{2}(t)=-6t,\ \lambda_{3}(t)=-2+\cos 4t$ and the 1st subsystem is unstable, and the 2nd and 3rd subsystems are stable. Then, by applying the Corollary 4, we can verify that $\sum_{i\in\mathbb{S}}\left(\mathbb{E}\left[\varphi_{i}(S_{i})\right]q_{i}+\sum_{j\in\mathbb{S}}q_{ij}\ln\mu_{ij}\right){\pi_{i}}<-1.17.$ Hence, system (26) with Markov switching signal is GUAS a.s. This example shows that our results still work for randomly switched systems with unstable subsystems.